OpenStudy (anonymous):

The quadratic 2x2 + 6x + 3 has what type of roots? A)Two Rational B)Two Complex C)One Rational Double D)Two Irrational The quadratic 2x2 + 6x + 3 has what type of roots? A)Two Rational B)One Rational Double C)Two Irrational D)Two Complex The quadratic 2x2 + 4x + 5 has what type of roots? A)Two Complex B)Two Irrational C)Two Rational D)One Rational Double

OpenStudy (anonymous):

b²-4ac>0 (has 2 root) b²-4ac<0 (has no root) b²-4ac=0 (has 1 root) ax² + bx + c =0

OpenStudy (anonymous):

do u understand wat i mean?

OpenStudy (anonymous):

No not really the choices sound confusing

OpenStudy (anonymous):

ok, for the first question, 2x² + 6x +3 is a function. We can write them in general form of quadrativ equation : ax² +bx +c = 0 2x²+6x+3=0 so u can see a = 2, b=6 c=3 input the value into this equation b²-4ac and see what is the value b²-4ac =6²-4(2)(3) =36-24>0 so this quadratic function has 2 different root. u can tryt he other 2 question

OpenStudy (anonymous):

So which choice would that be?

OpenStudy (anonymous):

i think it is a

jhonyy9 (jhonyy9):

2x2 +6x +3 =0 D=36-24=12 - so this signe that this equation will have 2 different real solutions

OpenStudy (anonymous):

Oh ok thank you